# Measuring multipartite quantum correlations by thermodynamic work extraction

### v1

@@ IV.1 Work deficit density for normal MPS under coarse-graining
In this subsection, we show that the work deficit density converges to the entanglement entropy under the coarse-graining when the system is described with the translationally invariant normal MPS with a fixed bond dimension $D_{B}$. We consider the situation that the $N$-site periodic system is partitioned into the $m$-site parties as $\{b_{n}\}_{n=1}^{L}$, where the relationship $L\equiv\frac{N}{m}$ holds. Under this partitioning, we can show the convergence of the work deficit density to the entanglement entropy as
$\delta_{\{b_{n}\}_{n=1}^{L}}=\frac{S_{\text{EE}}}{2}+\tilde{\mathcal{O}}(e^{-\frac{m}{2\xi}}),$ (33)
which is the main result of this section. Here, $\xi$ is the correlation length of this MPS, and $S_{\text{EE}}$ is the entanglement entropy for the subsystem much larger than $\xi$. The notation $\tilde{\mathcal{O}}(x)$ represents $\mathcal{O}(x\ln x)$. In the following, we provide a rough sketch of the proof of Eq. (33), while a precise discussion is presented in Supplemental Material.
What is crucial for proving this universal behavior is that normal MPS under sufficient coarse-graining converges to the following form Cirac et al. (2017, 2021); Piroli et al. (2021); Malz et al. (2024):
@@
where {|$\{\ket{i_{n}i_{n+1}}_{n}\}_{i_{n},i_{n+1}}$ denotes the orthonormal basis of party $b_{n}$. In this state, the bipartite-entangled state $\sum_{i_{n}}\sqrt{q_{i_{n}}}\ket{i_{n}}_{n}\ket{i_{n}}_{n+1}$ is shared between parties $b_{n}$ and $b_{n+1}$, which has a similar structure to the Bell-pair shared state in Fig. 5(b). Therefore, as discussed in Sec. II.3, work deficit densities of this state are reduced to the bipartite contribution, as
$\delta_{\{b_{n}\}_{n=1}^{L}}\simeq\frac{1}{L}\sum_{n=1}^{L}H(\{q_{i_{n}}\}),$ (35)
@@
where the optimal measurement protocol is to perform the projective measurement with basis {|$\{\ket{i_{n}i_{n+1}}_{n}\}_{i_{n},i_{n+1}}$ for all parties. Since the probability distributions $\{q_{i_{n}}\}$ in Eq. (35) are identical for all $n$ due to the translational invariance of the system, we can derive Eq. (33). The error term of $\tilde{\mathcal{O}}(e^{-\frac{m}{2\xi}})$ in Eq. (33) is the direct consequence of the error in the approximation of MPS in Eq. (34).
@@
IV.2 Detection of quantum phase transition with work deficit
@@
Figure 10: (a) Entanglement entropy for an MPS family defined in Eq. (30). Here, $S_{\rm EE}$ is defined as the entanglement entropy of a subsystem whose size much larger than the correlation length $\xi_{g}$, for each parameter $g$. Therefore, $S_{\rm EE}$ is ill-defined at $g=0$ due to the divergence of the correlation length. (b) Rough sketch of the behavior of work deficit density $\delta$ in the MPS family under coarse-graining. In the region $|g|\gg\frac{1}{m}$, $\delta$ converges to half of the entanglement entropy depicted in (a), while in the region $|g|\ll\frac{1}{m}$, $\delta$ is close to 00. Based on these analyses, we can conclude that $\delta$ undergoes drastic changes in the parameter region $|g|\sim\mathcal{O}(\frac{1}{m})$. It is important to note that this figure is a schematic derived from analytical calculations in the regions $|g|\ll\frac{1}{m}$ and $|g|\gg\frac{1}{m}$, and is not based on the calculations in the entire parameter range $-1\leq g\leq 1$.
@@
We apply the framework of coarse-graining to the MPS family defined in Eq. (30), to reveal its long-range quantum correlation structure. We consider the infinite system under periodic boundary condition, in the same way as Sec. III.3.3. Instead of conducting numerical calculations across the entire parameter range $-1\leq g\leq 1$ in a coarse-grained framework, we here perform the analysis solely within the regions $|g|\gg\frac{1}{m}$ and $|g|\ll\frac{1}{m}$ to grasp a rough sketch for the behavior of the work deficit density. Given that the correlation length (32) is approximately $\xi_{g}\simeq\frac{1}{2|g|}$ for $|g|\ll 1$, these regions correspond to the situations $\xi_{g}\ll m$ and $\xi_{g}\gg m$, respectively.
@@
In the region $|g|\gg\frac{1}{m}$, we can apply the discussion of Sec. IV.1, since $\xi_{g}\ll m$ is satisfied. Therefore, the work deficit density converges to the entanglement entropy, which can be calculated as
$S_{\rm EE}=\left\{\begin{aligned} &2\ln 2\ \ &(g<0)\\ &2h\left(\frac{1}{2}-\frac{\sqrt{g}}{1+g}\right)\ \ &(g>0)\end{aligned}\right.,$ (36)
where $h(p)$ denotes the binary entropy function. The behavior of entanglement entropy is also depicted in Fig. 9(a). We note that the size of the subsystem in calculating the entanglement entropy in Eq. (36) is taken much larger than the correlation length $\xi_{g}$.
@@
On the other hand, in the region $|g|\ll\frac{1}{m}$, work deficit density deviates from the entanglement entropy since the party size is smaller than the correlation length. Therefore, in this region, we take a different strategy to evaluate $\delta$. Here, we utilize the heuristic upper bound $U_{\Lambda^{*}}$ defined with the ansatz measurement protocol $\Lambda^{*}$ as
$\delta_{\{b_{n}\}_{n=1}^{L}}\leq U_{\Lambda^{*}}\leq\tilde{\mathcal{O}}(m|g|),$ (37)
@@
which showcases that work deficit density is close to 00 in this parameter region $|g|\ll\frac{1}{m}$. This behavior of work deficit density around $g=0$ can be understood as the consequence of the strong ferromagnetic order. For example, in the generalized GHZ state at $g=0$, the work deficit remains the same value $\Delta=\ln 2$ independent of $L$, resulting in $\delta=0$ in the infinite system $L\to\infty$.
@@
Figure 10(b) depicts the rough sketch of the behavior of $\delta$, reflecting these analytical calculations. It illustrates that the work deficit density converges to the entanglement entropy in the region $|g|\gg\frac{1}{m}$, while it rapidly approaches to zero in the region $|g|\ll\frac{1}{m}$. Therefore, by changing the size of the party $m$, the unsmoothness of the work deficit density around $g=0$ becomes sharper, showcasing the drastic change in the long-range structure of multipartite quantum correlation. On the other hand, the other kinks observed in Fig. 8(c), such as the one at $g=0.175$, shift or become dull under sufficient coarse-graining, since the work deficit converges to the entanglement entropy, which is smooth at these points as depicted in Fig. 10(a). This illuminates that these kinks are regarded as artifacts of optimization. We remark that Fig. 10(b) roughly predicts the behavior of the work deficit density based on the analysis in the regions $|g|\gg\frac{1}{m}$ and $|g|\ll\frac{1}{m}$. The behaviors in other parameter regions are merely interpolations from these extremes.
@@ V Multipartite one-way work deficit
In the foregoing sections, we focused on the multipartite work deficit in non-adaptive situation for simplicity. In this section, we discuss the work deficit with adaptive protocols. Specifically, we introduce the multipartite one-way work deficit, defined with the LOCC class called one-way closed LOCC, where fixed one-way classical communication and local adaptive measurements are allowed. Due to the restriction in communication direction, we can derive its expression exactly calculable in general few-body systems. Furthermore, we show that the efficient calculation method introduced in Sec. III is applicable also to this measure, thus significantly reducing numerical cost, while there remains an open question regarding the accuracy with this method.
@@ V.1 Definition of multipartite one-way work deficit
Here, $\mathcal{M}_{N}^{\to}$ denotes the set of measurement protocols whose operators are denoted as $\{\Pi_{y_{1}}\otimes\Pi_{y_{2}}^{y_{1}}\otimes\dots\Pi_{y_{N}}^{Y_{N-1}}\}_{Y_{N}}$, where $\{\Pi_{y_{n}}^{Y_{n-1}}\}_{y_{n}}$ is the set of the rank-one orthogonal projector on the $n$-th party depending on the measurement outcomes $Y_{n-1}$. The expression of Eq. (40) can be derived in essentially the same way as the derivation of $W_{l}$ in Eq. (4), which is explained in Appendix B.1.
@@
This expression shows that the multipartite one-way work deficit is calculated through the optimization of the local measurement bases $\{\Pi_{y_{1}}\},\{\Pi_{y_{2}}^{y_{1}}\},\dots\{\Pi_{y_{N}}^{Y_{N-1}}\}$, which can be performed exactly for general few-body systems. One can also show that Eq. (41) can be rewritten using the quantum relative entropy as
@@ V.2 Efficient calculation method for multipartite one-way work deficit
We show that the efficient calculation method introduced in Sec. III is also applicable to the one-way work deficit, although there remains an open question in the error bound. Similarly to Sec. III, we define the one-way work deficit density as
@@
where $N$ denotes the number of parties. This represents the leading term of $\Delta^{\to}$ in many-body systems, and the numerical cost for its direct calculation explodes in $e^{e^{\mathcal{O}(N)}}$, as evident from Eq. (41). In the efficient calculation method, we estimate $\delta^{\to}$ by utilizing efficiently computable lower and upper bounds:
@@
For systems described with the MPS, the numerical costs for these bounds are proportional to the system size. Moreover, for translationally invariant infinite systems described with normal MPS, these bounds can be calculated with constant numerical cost.
The validity of these bounds can be demonstrated through calculations in some examples. For instance, in the AKLT state and the cluster state of a one-dimensional infinite system, we can exactly calculate the one-way work deficit density using these bounds as follows:
@@
The calculation procedures follow those for $\delta$ in Secs. III.3.1 and III.3.2, with further details provided in Supplemental Material.
@@ VI Summary and discussion
Importantly, we introduced an efficient calculation method for this measure in many-body systems described with MPS, in Sec. III.1 and III.2. While the numerical cost for directly calculating the multipartite work deficit grows exponentially with system size $N$, the cost for our efficient method is linear with respect to $N$, enabling significant cost reduction. This method is applicable not only to the multipartite work deficit but also to another measure called global quantum discord Rulli and Sarandy (2011), which highlights its potential applicability to various measures. In Sec. III.3, we demonstrated this method in several MPS examples, including the AKLT state, the cluster state, and an MPS family Wolf et al. (2006). We also introduced the framework of coarse-graining of the multipartite work deficit in Sec. IV. By applying coarse-graining to the MPS family, we observed the drastic change of multipartite quantum correlaiton structure at the quantum phase transition point.
@@
A remaining issue is to investigate whether our efficient calculation method for MPS can be applied to broader classes of many-body states. Such extension may be straightforward for the systems described with projected-entangled pair states (PEPS), matrix product operators (MPO), and projected entangled pair operators (PEPO), which are the generalizations of MPS to higher-dimensional and/or mixed states Cirac et al. (2021, 2017). In fact, these states can be represented with fixed bond dimensions independent of the system size, which is essential for our calculation method. On the other hand, extension to states that exhibit volume-law entanglement entropy is much more nontrivial, since the bond dimension for these states grows exponentially with the system size. It is an important open question to develop novel approaches for evaluating multipartite work deficit in such states.
@@
Another remaining issue is to study the multipartite work deficit defined with fully adaptive LOCC. In this work, we introduced the work deficits defined with the LOCC classes called zero-way and one-way closed LOCC, where the set of allowed operations are reasonably restricted. This restriction is crucial for deriving their closed formula which can be calculated exactly for general few-body states, while maintaining clear operational meanings based on work extraction. On the other hand, the investigation of the measure defined with fully adaptive LOCC, in which the adaptive local operations and arbitrary classical communication are allowed, poses a greater challenge due to the extensive range of allowable LOCC protocols. Since studies on such cases have so far been limited to demonstrations in a few examples Oppenheim et al. (2002); Horodecki et al. (2005), it remains a future task how to calculate the work deficit under the fully adaptive situation in both few-body and many-body systems.
@@
It is also important to explore how the multipartite work deficit reveals the structures of quantum correlations in quantum many-body systems. While previous works have studied quantum correlations at zero and finite temperature many-body systems Alicki et al. (2010); Hastings (2011); Kuwahara and Saito (2022); Wolf et al. (2008); Kuwahara et al. (2021); Anshu et al. (2020); Cirac et al. (2011); Schuch et al. (2013); Kato and Brandão (2019); Anshu et al. (2022); Kuwahara et al. (2020); Kato and Brandao (2019), they primarily focused on bipartite or tripartite measures, leaving the structure of multipartite quantum correlation largely unexplored. The multipartite work deficit introduced in this work would serve as a probe for elucidating such multipartite structures. Particularly, its convergence to entanglement entropy in normal MPS, as discussed in Sec. IV.1, represents a significant step forward. Clarifying the conditions under which this convergence to bipartite entanglement occurs poses an open question.
@@
Finally, we discuss the relevance of our work to experiments. Recently, single-site measurement and manipulation of quantum many-body systems Bakr et al. (2009); Weitenberg et al. (2011), as well as thermodynamic work extraction from quantum systems Cottet et al. (2017); Masuyama et al. (2018); Naghiloo et al. (2018), are becoming feasible on real experimental platforms, such as ultracold atoms and superconducting qubits. The multipartite work deficit has fully operational meaning, as it is defined based on the work extraction through site-wise operations, which would be feasible with current or near-future quantum technologies. Therefore, the multipartite work deficit is expected to serve as an experimental probe of multipartite structure of quantum correlations in many-body quantum systems. It would also be an interesting perspective to investigate the role of quantum correlations in thermodynamics in such experimental platforms.

### v2

@@ IV.1 Work deficit density for normal MPS under coarse-graining
In this subsection, we show that the work deficit density $\delta_{m}$ converges to half of the entanglement entropy by increasing number of sites per party $m$. Especially, we focus on the translationally invariant normal MPS Cirac et al. (2017, 2021) under periodic boundary condition. We consider an $N$-site system divided into $m$-site parties, where the number of parties is denoted as $M\equiv\frac{N}{m}$. In this setup, we can derive the following equality, which is the main result of this section:
$\delta_{m}=\frac{S_{\text{EE}}}{2}+\tilde{\mathcal{O}}(e^{-\frac{m}{2\xi}}),$ (33)
where $\xi$ is the correlation length of this MPS, $S_{\text{EE}}$ is the entanglement entropy for subsystem much larger than $\xi$, and the notation $\tilde{\mathcal{O}}(x)$ represents $\mathcal{O}(x\ln x)$. In the following, we provide an intuitive discussion of why Eq. (33) is satisfied, while the rigorous proof is presented in Supplemental Material as Theorem 4.
The essence of the derivation of Eq. (33) is the fact that under coarse-graining, a normal MPS converges to the state where the bipartite-entangled pairs are shared between the neighbouring parties, as shown in Fig. 9(b). This convergence is the consequence of the renormalization group transformation, as discussed e.g., in Refs. Cirac et al. (2017, 2021); Piroli et al. (2021); Malz et al. (2024). The state in Fig. 9(b) is represented as
@@
where {|$\{\ket{jk}_{n}\}_{j,k=1}^{D_{B}}$ denotes the orthonormal basis of $n$-th party $b_{n}$. In this representation, each $m$-site party is effectively regarded as a $D_{B}^{2}$-dimensional system due to the fixed bond dimension of |$\ket{\psi}$. Since Eq. (34) is a generalized version of the Bell-pair-shared state in Fig. 5(b), where the Bell pair is replaced with the bipartite-entangled pair $\sum_{i}\sqrt{q_{i}}\ket{i}_{n}\ket{i}_{n+1}$, the work deficit density for this state is reduced to the bipartite contribution as follows:
$\delta_{m}\simeq\frac{1}{M}\sum_{n=1}^{M}H(\{q_{i}\})=H(\{q_{i}\}).$ (35)
@@
Here, the optimal measurement protocol for $\delta_{m}$ is to perform the projective measurement with basis {|$\{\ket{jk}_{n}\}_{j,k}$ for all parties. Since $S_{\text{EE}}=2H(\{q_{i}\})$ as shown in Fig. 9(b), we can derive Eq. (33). The error term $\tilde{\mathcal{O}}(e^{-\frac{m}{2\xi}})$ in Eq. (33) is the consequence of the error in the approximation in Eq. (34).
@@
IV.2 Detection of quantum phase transition with work deficit density
@@
Figure 10: (a) Entanglement entropy for the MPS family (Eq. (30)). Here, $S_{\rm EE}$ is defined as the entanglement entropy of a subsystem whose size is much larger than the correlation length $\xi_{g}$, for each parameter $g$. Therefore, $S_{\rm EE}$ is ill-defined at $g=0$ due to the divergence of the correlation length. (b) Rough sketch of the behavior of work deficit density $\delta_{m}$ in the MPS family under coarse-graining. In the region $|g|\gg\frac{1}{m}$, $\delta_{m}$ converges to half of the entanglement entropy depicted in (a), while in the region $|g|\ll\frac{1}{m}$, $\delta_{m}$ is close to zero. Based on these analyses, we can conclude that $\delta_{m}$ undergoes drastic changes in the parameter region $|g|\sim\mathcal{O}(m^{-1})$. It is important to note that this figure is a rough sketch derived from analytical calculations in the regions $|g|\ll\frac{1}{m}$ and $|g|\gg\frac{1}{m}$, and is not based on the calculations in the entire parameter range $-1\leq g\leq 1$.
@@
We apply the framework of coarse-graining to the MPS family defined in Eq. (30), to reveal its long-range quantum correlation structure. We consider the system under periodic boundary condition in the thermodynamic limit, in the same way as Sec. III.3.3. Instead of conducting numerical calculations across the entire parameter range $-1\leq g\leq 1$ in a coarse-grained framework, we here perform the analysis solely within the regions $|g|\gg\frac{1}{m}$ and $|g|\ll\frac{1}{m}$ to grasp a rough sketch for the behavior of the work deficit density. Given that the correlation length (32) is approximately $\xi_{g}\simeq\frac{1}{2|g|}$ for $|g|\ll 1$, the regions $|g|\gg\frac{1}{m}$ and $|g|\ll\frac{1}{m}$ correspond to the situations of $\xi_{g}\ll m$ and $\xi_{g}\gg m$, respectively.
@@
In the region $|g|\gg\frac{1}{m}$, we can utilize Eq. (33), since $\xi_{g}\ll m$ is satisfied. Therefore, the work deficit density converges to half of the entanglement entropy of subsystem much larger than the correlation length $\xi_{g}$, which can be calculated as follows:
$\frac{S_{\text{EE}}}{2}=\left\{\begin{aligned} &\ln 2\ \ &(g<0)\\ &h\left(\frac{1}{2}-\frac{\sqrt{g}}{1+g}\right)\ \ &(g>0)\end{aligned}\right.,$ (36)
where $h(p)\equiv-p\ln p-(1-p)\ln(1-p)$ denotes the binary entropy function. It is depicted in Fig. 10(a).
@@
On the other hand, in the region $|g|\ll\frac{1}{m}$, work deficit density deviates from the entanglement entropy since the party size $m$ is smaller than the correlation length $\xi_{g}$. Therefore, in this region, we take a different strategy to evaluate $\delta_{m}$. To be specific, we can derive the heuristic upper bound
$\delta_{m}\leq\tilde{\mathcal{O}}(m|g|),$ (37)
@@
by utilizing a certain ansatz measurement protocol SM . From this inequality, we can see that $\delta_{m}$ is close to zero in the region $|g|\ll\frac{1}{m}$. Such a behavior of work deficit density can be understood as the consequence of the long-range correlation around $g=0$. For example, in the generalized GHZ state at $g=0$, the total work deficit remains the same value $\ln 2$ independent of the number of parties $M$, resulting in $\delta_{m}=0$ in the thermodynamic limit $M\to\infty$.
@@
Figure 10(b) shows the rough sketch of the behavior of work deficit density, reflecting these analytical calculations. This figure illustrates that $\delta_{m}$ converges to half of the entanglement entropy in the region $|g|\gg\frac{1}{m}$, while it rapidly approaches to zero in the region $|g|\ll\frac{1}{m}$. Therefore, by changing the size of the party $m$, the unsmoothness of the work deficit density around $g=0$ becomes sharper, showcasing the drastic change in the long-range structure of multipartite quantum correlation. On the other hand, the other kinks observed in Fig. 8(c), such as the one at $g=0.175$, shift or become dull under sufficient coarse-graining, since $\delta_{m}$ converges to $\frac{S_{\rm EE}}{2}$, which is smooth at these points as depicted in Fig. 10(a). This illuminates that these kinks are regarded as artifacts of optimization. We remark that Fig. 10(b) roughly predicts the behavior of the work deficit density based on the analysis in the regions $|g|\gg\frac{1}{m}$ and $|g|\ll\frac{1}{m}$. The behaviors in other parameter regions are merely interpolations from these extremes.
@@ V Multipartite one-way work deficit
In the foregoing sections, we focused on the multipartite work deficit in non-adaptive situation for simplicity. In this section, we discuss the work deficit with adaptive protocols. Specifically, we introduce the multipartite one-way work deficit, defined with the LOCC class called one-way closed LOCC, where fixed one-way classical communication and local adaptive measurements are allowed. Due to the restriction in communication direction, we can derive its expression exactly calculable in general few-body systems. Furthermore, we show that the efficient calculation method introduced in Sec. III is applicable also to this measure, thus significantly reducing numerical cost, while there remains an open question regarding the accuracy of this method.
@@ V.1 Definition of multipartite one-way work deficit
Here, $\mathcal{M}_{N}^{\to}$ denotes the set of measurement protocols whose operators are denoted as $\{\Pi_{y_{1}}\otimes\Pi_{y_{2}}^{y_{1}}\otimes\dots\Pi_{y_{N}}^{Y_{N-1}}\}_{Y_{N}}$, where $\{\Pi_{y_{n}}^{Y_{n-1}}\}_{y_{n}}$ is the set of the rank-one orthogonal projectors on the $n$-th party depending on the measurement outcomes $Y_{n-1}$. The derivation of $W_{l}^{\to}$ in Eq. (40) is essentially the same as that of $W_{l}$ in Eq. (4), which is explained in detail in Appendix B.
@@
This expression shows that the multipartite one-way work deficit is calculated through the optimization of the local measurement bases $\{\Pi_{y_{1}}\},\{\Pi_{y_{2}}^{y_{1}}\},\dots\{\Pi_{y_{N}}^{Y_{N-1}}\}$, which can be performed exactly for general few-body systems. One can also show that Eq. (41) is rewritten using the quantum relative entropy as
@@ V.2 Efficient calculation method for multipartite one-way work deficit
We show that the efficient calculation method introduced in Sec. III is also applicable to the one-way work deficit, although there remains an open question in the error bound. As the leading term of $\Delta^{\to}$ in many-body systems, we define the one-way work deficit density as
@@
where $N$ denotes the number of parties. The numerical cost for directly calculating $\delta^{\to}$ explodes in $e^{e^{\mathcal{O}(N)}}$, since we need to optimize all the measurement bases $\{\Pi_{y_{1}}\}_{y_{1}},\{\Pi_{y_{2}}^{y_{1}}\}_{y_{1},y_{2}},\dots\{\Pi_{y_{N}}^{Y_{N-1}}\}_{Y_{N}}$, as shown Eq. (41). In the efficient calculation method, we evaluate $\delta^{\to}$ by utilizing efficiently computable lower and upper bounds:
@@
For systems described with MPS, the numerical costs for these bounds $L_{l}^{\to}$ and $U_{l}^{\to}$ are proportional to the system size. Moreover, for translationally invariant normal MPS in the thermodynamic limit, these bounds can be calculated with constant numerical cost.
The validity of these bounds can be demonstrated through calculations in some examples. For instance, in the AKLT state and the cluster state in the thermodynamic limit, we can exactly calculate the one-way work deficit density using these bounds as follows SM :
@@ VI Summary and discussion
Importantly, we introduced an efficient calculation method for this measure in many-body systems described with MPS, in Secs. III.1 and III.2. While the numerical cost for directly calculating the multipartite work deficit grows exponentially with system size $N$, the cost for our efficient method is linear with respect to $N$, enabling significant cost reduction (Table 1). This method is applicable not only to the multipartite work deficit but also to another measure called global quantum discord Rulli and Sarandy (2011), which highlights its potential applicability to various measures. In Sec. III.3, we demonstrated this method in several MPS examples, including the AKLT state, the cluster state, and an MPS family Wolf et al. (2006). We also introduced the framework of coarse-graining of the multipartite work deficit in Sec. IV. By utilizing this framework in the MPS family, we observed the drastic change of multipartite quantum correlaiton structure at the quantum phase transition point.
@@
A remaining issue of this work is to investigate whether our efficient calculation method for MPS can be applied to broader classes of many-body states. Such extension may be straightforward for the systems described with projected-entangled pair states (PEPS), matrix product operators (MPO), and projected entangled pair operators (PEPO), which are the generalizations of MPS to higher-dimensional and/or mixed states Cirac et al. (2021, 2017). In fact, these states can be represented with fixed bond dimensions independent of the system size, which is essential for our calculation method. On the other hand, extension to states that exhibit volume-law entanglement entropy is much more nontrivial, since the bond dimension for these states grows exponentially with the system size. It is an important open question to develop novel approaches for evaluating multipartite work deficit in such states.
@@
Another future task is to study the multipartite work deficit defined with fully adaptive LOCC. In this work, we introduced the work deficits defined with the LOCC classes called zero-way and one-way closed LOCC, where the set of allowed operations are reasonably restricted. This restriction is crucial for deriving their closed formula which can be calculated exactly for general few-body states, while maintaining clear operational meanings based on work extraction. On the other hand, the investigation of the measure defined with fully adaptive LOCC, in which the adaptive local operations and arbitrary classical communication are allowed, poses a greater challenge due to the extensive range of allowable protocols. Since studies on such a case have so far been limited to demonstrations in a few examples Oppenheim et al. (2002); Horodecki et al. (2005), it remains a future task to establish calculation methods for general few-body and many-body systems.
@@
It is also important to explore how the multipartite work deficit reveals the structures of quantum correlations in quantum many-body systems. While the quantum correlations of many-body systems at zero and finite temperature have been extensively studied in the previous works Alicki et al. (2010); Kuwahara and Saito (2022); Wolf et al. (2008); Kuwahara et al. (2021); Anshu et al. (2020); Cirac et al. (2011); Schuch et al. (2013); Kato and Brandão (2019a); Anshu et al. (2022); Kuwahara et al. (2020); Kato and Brandão (2019b), they primarily focused on bipartite or tripartite measures, leaving the structure of multipartite quantum correlation largely unexplored. We expect that the multipartite work deficit introduced in this work would serve as a probe for elucidating such multipartite structures. Particularly, its convergence to entanglement entropy in normal MPS, as discussed in Sec. IV.1, represents a significant step forward. Clarifying the conditions under which this convergence to bipartite entanglement occurs poses an open question.
@@
Finally, we discuss the relevance of our work to experiments. Recently, single-site measurement and manipulation of quantum many-body systems Bakr et al. (2009); Weitenberg et al. (2011), as well as thermodynamic work extraction from quantum systems Cottet et al. (2017); Masuyama et al. (2018); Naghiloo et al. (2018), are becoming possible on real experimental platforms, such as ultracold atoms and superconducting qubits. The multipartite work deficit has fully operational meaning, as it is defined based on the work extraction through site-wise operations, which would be feasible with these current or near-future quantum technologies. Therefore, the multipartite work deficit is expected to serve as an experimental probe of multipartite structure of quantum correlations in many-body quantum systems. It would also be an interesting perspective to investigate the role of quantum correlations in thermodynamics in such experimental platforms.